Solve the recurrence relation of an=3an-3an-2-an-3 n>=3 a0=1 a1=-2 a2=-1
1
Expert's answer
2022-04-07T13:25:30-0400

"a_n-3a_n+3a_{n-2}+a_{n-3}=0"

"-2a_n+3a_{n-2}+a_{n-3}=0"

Characteristics equalation

"-2x^3+3x+1=0"

"Q=(a^2-3b)\/9=0.5"

"R=(2a^3-9ab+27c)\/54=-0.25"

"S=Q^3-R^2=0.06" >0

"\phi=1/3 arccos(R/\sqrt{Q^3})"

"x_1=-2\sqrt{Q} cos(\phi)-a/3=-1"

"x_2=-2\sqrt{Q} cos(\phi+2/3\pi)-a/3=1.37"


"x_3=-2\sqrt{Q} cos(\phi-2/3\pi)-a/3=-0.37"

"a_n=a(-1)^n+(1.37)^nb+(-0.37)^nc"

"a_0=a+b+c=1"

"a_1=-a+1.37b-0.37c=-2"

"a_2=a+1.8769b+0.1369c=-1"

"a_0+a_1=2.37b+0.63c=-1"

"a_1+a_2=3.2469b-0.2331c=-3"

"b=(-1-0.63c)\/2.37"

"3.2469(-1-0.63c)-0.2331c=-3"

"-3.2469-2.045547c-0.2431c=-3"

"-2.288647c=0.2469"

"c=-0.0108"

"b=-0.419"

"a=1-b-c=1+0.0108+0.419=1.43"

"a_n=1.43(-1)^n-0.419(1.37)^n-0.0108(-0.37)^n"